40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2242     Entry time: Thu Sep 27 10:58:58 2018
Author: awade and anchal 
Type: DailyProgress 
Category: FSS 
Subject: First FSS lock 36 MHz north path 

We hooked up the 36 MHz crystal source to the TTFSS boxes and EOM driver yesterday and scanned the north cavity to see the error signal.  We got a bunch of noisy junk around each resonance rather than a clean PDH function. This post is diagnosing what the issue is.

Some setup details

The the frequency source is a preamplified OCXO crystal sources (see PSL:2235, for info and links) that outputs about +17 dBm out of the LO SMA and +26 dBm out of the EOM driver SMA. 

We attenuated the LO output with a 10 dB minicircuits attenuator (VAT-10+) to bring the power down to +7dBm into the TTFSS field box.  The TTFSS field box RF board (LIGO-D0901894) then amplifies the signal up to +17 dBm to be used in a level 17 mixer for demodulating the photodetector signal. 

The EOM drivers need about -3 dBm to produce a modulation depth of about 0.3.  The EOM output of the OXCO box was attenuated with a 9 dB attenuator (VAT-9+) followed by 20 dB attenuator (VAT-20+) to give a measured -3 dBm at 36 MHz (north).  

The new photodetector RF transfer functions are detailed in PSL:2240 and the tuned EOM drivers TF are posted in PSL:2236.   All these units are now powered from the same ±18 V power strip, eliminating a bunch of independent  bench top power supplies that were previously being used.  The benefit of this is now they all share a common ground referenced to the same central grounding point at the rack (in a tree like network) and that we are removing flaky banana plug connections everywhere in the experiment (bananas are bad PSL:1705, PSL:1759, PSL:1760, PSL:1843, 40m:505240m:5098).  Previously many elements in the FSS electronics were not grounded properly.  

Testing RF demodulation

A 37 MHz +7 dBm sine wave was injected into the TTFSS field box LO port and a +0 dBm signal was injected at 37.000010 MHz (10 Hz offset) to see if we were actually getting a signal.  We got a 656 mV peak to peak sine wave out of the mixer monitor port of the FSS box (50 Ω terminated). From the internals of the TTFSS box (see LIGO- D0901894) the ADT1-1 transformer couplers have an insertion loss of 0.31 dB @ 40 MHz and the level-17 JMS-1H has a conversion loss of order 5.90 dB (middle of its band).  The RF notched filter has a pass band insertion loss of 0.45 dB (see previous PSL post).  Before reading out at the monitor port the signal is amplified by U3 on LIGO-D040105 by 3.16 (+ 5.0 dB).  This should give a final Vpp of 0.978 V.  So we are seeing less power conversion than we expected. 

It should be noted that the LO path in the FSS board has a +10 dB amplifier followed by a ADC-10-4 coupler with an insertion loss of order 0.74 dB.  It could be that the mixer is not seeing the full +17 dBm and this is why the conversion looks a little low.

Other than this slightly lower than expected conversion this RF demodulation stage seems to be working as expected.

EOM driver saturating

The EOM driver was receiving -3 dBm of power.   However as it turns out with the x 155 gain the output measured at the RF monitor port of the EOM driver was order 440 mV, or 1.02 rad which is way too much for now.  The attenuation was adjusted so the power into the EOM driver was  -6 dBm, giving 310 mVpp at the monitor port.

A ground loop issue with the North laser controler

I plugged the RF output of the PMC reflection PD into a spectrum analyzer.  As we scan accross the resonance I see a pair of 500 kHz peaks either side of the fundamental mode.  When I lock the PMC and look at the RF out of the refcav reflection PD I see the same pair of 437 kHz sidebands around the 21.5 MHz peak.  I see a similar thing close to DC with a  447 kHz line and 877 kHz.  These are the same features mixed up to 21.5 MHz.  

When I disconnect the PZT in of the north laser controler these peaks go away.  It looks like we have some kind of ground loop issue.  

 

ELOG V3.1.3-